# suppSupplementary References

### v2

@@ 1 Introduction
[…]
@@
for $p\in\mathbb{N}$. Here, $F^{-1}$ and $G^{-1}$ denote the generalized quantile functions, given by $F^{-1}(\tau)=\inf\{x\in\mathbb{R}\mid\tau\leq F(x)\}$, $\tau\in[0,1]$ and accordingly for $G^{-1}$, i.e., the left-inverses of the CDFs. The Cramér distance arises as the square of the special case $p=2$ within the class of $l_{p}$ distances,
@@
While extensive work has been done on decomposing proper scoring rules (Hersbach, 2000; Bröcker, 2009; Dimitriadis et al., 2021; Bracher et al., 2021, among many others), the literature on decompositions of divergence functions into interpretable components is sparse. An exception is the exact decomposition of the 2-Wasserstein distance into the squared differences of the distributions’ means and standard deviations, together with an analytically known remainder term capturing differences in shape (del Barrio et al.,, 1999; Irpino and Verde,, 2015), which has recently been used in applications by Schefzik et al., (2021) and Lorenzo and Arroyo, (2022). In contrast to this moment-based approach, which compares summary statistics that arise for each of the distributions separately, our decompositions are fully nonparametric in the sense of aggregating (integrating over) all distributional differences of $F$ and $G$. Furthermore, our approach does not require a remainder term and is applicable to a range of divergence functions.
@@
The supplementary text contains proofs and derivations together with additional illustrations, counterexamples and other details. Replication code is available at https://github.com/resinj/replication_SD-Decomp.
@@ 2.1 The area validation metric
$\operatorname{AVM}_{\alpha}(F,G)=\,\underbrace{\operatorname{Shift}_{\alpha,+}^{\operatorname{AVM}}(F,G)}_{\text{``}F\text{ shifted up''}}\,+\,\underbrace{\operatorname{Shift}_{\alpha,-}^{\operatorname{AVM}}(F,G)}_{\text{``}F\text{ shifted down''}}\,+\,\underbrace{\operatorname{Disp}_{\alpha,+}^{\operatorname{AVM}}(F,G)}_{\text{``}F\text{ more dispersed''}}\,+\,\underbrace{\operatorname{Disp}_{\alpha,-}^{\operatorname{AVM}}(F,G)}_{\text{``}F\text{ less dispersed''}}.$ (6)
Using $[z]_{+}:=\max(z,0)$ for the positive part of a real number $z\in\mathbb{R}$, we define the $\alpha$-wise components
@@
We henceforth refer to the two components with subscript ‘$+$’ as the plus components as they quantify how $F$ is, relative to $G$, shifted upwards and has an increased dispersion, respectively. Similarly, we refer to the terms with subscript ‘--’ as the minus components.
The decomposition terms in (7)–(8) attribute the overall difference between the central $\alpha$-interval endpoints of $F$ and $G$—which is captured by the integrand $\operatorname{AVM}_{\alpha}(F,G)$ in (5)—to the components by more intricate interval comparisons. In a nutshell, the shift components capture twice the distance that one of the $\alpha$-intervals needs to be moved to lie within the other, while the dispersion components measure by how much one needs to (de)compress the $\alpha$-interval of $G$ to have the same length as the $\alpha$-interval of $F$.
@@
If, on the other hand, the $G$-interval is wider than the $F$-interval, the difference in interval length is attributed to the minus dispersion component. Analogously, if the intervals are ordered differently, their difference in location is attributed to the minus shift component. On a technical level, these separations into plus and minus terms are achieved by the $[\cdot]_{+}$ operator in (7) and (8), respectively.
@@ 2.2 The $p$-Wasserstein distance
We next present a generalized decomposition for the $p$-th power of the $p$-Wasserstein distance in (2) for $p\in\mathbb{N}$. Here, we directly state the decomposition in its integral-form and dispense with a disaggregated $\alpha$-wise treatment as in (6), which can be obtained by simply omitting the integrals in the following formulas.
@@ 2.3 The Cramér distance
[…]
@@ 3.1 Basic properties
For any $p\in\mathbb{N}$ and $\operatorname{D}\in\{\operatorname{AVM},\operatorname{WD}_{p},\operatorname{CD}\}$, we have
@@
Hence, the properties that are outlined for the plus components (with subscript ‘$+$’) also apply to the minus counterparts (with subscript ‘--’) by symmetry.
@@
Let $F$ and $G$ be symmetric distributions with central medians $m_{F}$ and $m_{G}$ and $\operatorname{D}\in\{\operatorname{AVM},\operatorname{WD}_{p},\operatorname{CD}\}$, $p\in\mathbb{N}$.
@@
    For any $\operatorname{D}\in\{\operatorname{AVM},\operatorname{WD}_{p},\operatorname{CD}\}$ with $p\in\mathbb{N}$, it holds that
@@
    The central medians of $F$ and $G$ are $m_{F}=s_{F}m_{H}+\ell_{F}$ and $m_{G}=s_{G}m_{H}+\ell_{G}$, respectively, where $m_{H}$ denotes the central median of $H$. For any $p\in\mathbb{N}$, if $m_{F}\leq m_{G}$, then $\operatorname{Shift}_{+}^{\operatorname{WD}_{p}}(F,G)=0$. Moreover, if the median of $H$ is unique,
@@
We finally consider the case of normal distributions, which, arguably, form the most prominent location-scale family. Let $F=\mathcal{N}(\mu_{F},\sigma_{F}^{2})$ and $G=\mathcal{N}(\mu_{G},\sigma_{G}^{2})$ be normal distributions with means $\mu_{F}$ and $\mu_{G}$ and variances $\sigma_{F}^{2}$ and $\sigma_{G}^{2}$, respectively. In what follows, we use the shorthand notations $\widetilde{\mu}=|\mu_{F}-\mu_{G}|$, $\widetilde{\sigma}=|\sigma_{F}-\sigma_{G}|$ and $\widetilde{\rho}=\sqrt{\sigma_{F}^{2}+\sigma_{G}^{2}}$. We further denote the standard normal density and distribution function by $\phi(\cdot)$ and $\Phi(\cdot)$, respectively. Here, we provide closed-form expressions for the decompositions of the $\operatorname{AVM}$ and the $\operatorname{CD}$. More involved formulas for the $\operatorname{WD}_{p}$ with arbitrary $p\in\mathbb{N}$ are given in Supplement S2.1.
@@
$\operatorname{AVM}(F,G)$ $=\widetilde{\mu}\big{(}2\Phi(\widetilde{\mu}/\widetilde{\sigma})-1\big{)}+2\widetilde{\sigma}\phi(\widetilde{\mu}/\widetilde{\sigma}).$
@@
$\operatorname{Shift}_{+}^{\operatorname{CD}}(F,G)=\widetilde{\mu}\big{(}1-2\Phi(-\widetilde{\mu}/\widetilde{\rho})\big{)}-2\tau\big{(}\phi(0)-\phi(\widetilde{\mu}/\widetilde{\rho})\big{)}.$
@@ 3.2 Agreement and differences across distance measures
For any two distributions $F$ and $G$, and $p\in\mathbb{N}$, it holds that
@@
For any two distributions $F$ and $G$, and $p\in\mathbb{N}$, it holds that
@@
Let $F$ and $G$ be symmetric distributions, $F\neq G$, and $q>p\geq 1$ be positive integers. Then, it holds that
@@ 4 Compatibility with stochastic order relations
[…]
@@
[…]
@@ 4.1 Connections to the dispersive order
The distribution $F$ is said to be larger than $G$ in dispersive order, $F\geq_{\text{D}}G$, if for all $10<\xi<\tau<$ 1,
@@
The dispersive order turns out to be too strong for a logical equivalence with zero dispersion components to hold. To establish such an equivalence, we consider a weaker order relation. The distribution $F$ is said to be larger than $G$ in weak dispersive order, $F\geq_{\mathrm{wD}}G$, if for all $10.5<\tau<$ 1,
@@
Let $F$ and $G$ be probability distributions with continuous quantile functions $F^{-1}$ and $G^{-1}$, respectively. Then, for any $\operatorname{D}\in\{\operatorname{AVM},\operatorname{WD}_{p},\operatorname{CD}\}$, $p\in\mathbb{N}$, we have that
@@
Subject to the regularity condition of continuous quantile functions, which is further discussed below, part (a) of the theorem establishes an equivalence between a weak dispersive ordering of the distributions and a zero dispersion component in our decompositions. As shown in part (b), the equivalence extends naturally to an equivalence between a corresponding strict ordering and a unique nonzero dispersion component. Notably, the equivalences and implications in Theorem 4.1 hold for all considered divergence measures $\operatorname{D}\in\{\operatorname{AVM},\operatorname{WD}_{p},\operatorname{CD}\}$, $p\in\mathbb{N}$, by the equivalences from Proposition 3.7.
@@
The issue encountered in Example 4.3 could be avoided by slightly adapting the definition of the weak dispersive order in (24) to require for any $10.5\leq\tau<$ 1 that
@@ 4.2 Connections to the usual stochastic order
Figure 8: Connections between the stochastic order relations introduced in Section 4.2 and the shift components of the $\operatorname{CD}$ and $\operatorname{WD}_{p}$ divergence measures (for any $p\in\mathbb{N}$).
@@
The distribution $F$ is said to be larger than $G$ in the usual stochastic order, $F\geq_{\mathrm{S}}G$, if $F(x)\leq G(x)$ for all $x\in\mathbb{R}$, or, equivalently, if $F^{-1}(\tau)\geq G^{-1}(\tau)$ for all $10<\tau<$ 1. The usual stochastic order is a partial order, and its definition can be reformulated as
[…]
for all $10.5<\tau<$ 1. The term used for this unconventional characterization arises in the shift components (see (7) with $\tau=\frac{1+\alpha}{2}$), thereby serving as a natural starting point to investigate the connection.

### v3

@@ 1 Introduction
Here, the shift components $\operatorname{Shift}_{\pm}^{\operatorname{D}}$ with $\pm\in\{+,-\}$ quantify differences in location, while the dispersion components $\operatorname{Disp}_{\pm}^{\operatorname{D}}$ measure differences in variability between $F$ and $G$. The signed components (with subscript ‘++’ and ‘-’) attribute parts of the distance to upwards and downwards shifts, and more or less dispersion of $F$ relative to $G$. Of course, the components in (1) are functions of the pair of distributions $(F,G)$. For the sake of brevity, however, we will sometimes omit this dependence and use the shorthands $\operatorname{Shift}_{\pm}^{\operatorname{D}}=\operatorname{Shift}_{\pm}^{\operatorname{D}}(F,G)$ and $\operatorname{Disp}_{\pm}^{\operatorname{D}}=\operatorname{Disp}_{\pm}^{\operatorname{D}}(F,G)$. These refer to the components of $\operatorname{D}(F,G)$ between for two generic distributions $F$ and $G$ or a pair of distributions which becomes clear from the context.
@@
for $p\in[1,\infty)$. Here, $F^{-1}$ and $G^{-1}$ denote the quantile functions, given by $F^{-1}(\tau)=\inf\{x\in\mathbb{R}\mid\tau\leq F(x)\}$, $\tau\in[0,1]$ and accordingly for $G^{-1}$, i.e., the left-inverses of the CDFs. The Cramér distance arises as the square of the special case $p=2$ within the class of $l_{p}$ distances,
@@
While extensive work has been done on decomposing proper scoring rules (Hersbach, 2000; Bröcker, 2009; Dimitriadis et al., 2021; Bracher et al., 2021, among many others), the literature on decompositions of divergence functions into interpretable components is sparse. An exception is the exact decomposition of the 2-Wasserstein distance into the squared differences of the distributions’ means and standard deviations, together with an analytically known remainder term capturing differences in shape (del Barrio et al.,, 1999; Irpino and Verde,, 2015), which has recently been used in applications by Schefzik et al., (2021) and Lorenzo and Arroyo, (2022). Remark 2.4 illustrates that in contrast to this moment-based approach, which compares summary statistics that arise for each of the distributions separately, our decompositions are fully nonparametric in the sense of aggregating (integrating over) all distributional differences of $F$ and $G$. Furthermore, our approach does not require a remainder term and is applicable to a range of divergence functions.
@@
The supplementary text contains proofs and derivations together with details on how the decompositions can be computed or approximated in practice, additional illustrations, counterexamples and further details. Replication code is available at https://github.com/resinj/replication_SD-Decomp.
@@ 2.1 The area validation metric
[…]
Using $z]_{+}:=\max(z,0)$ for the positive part of a real number $z\in\mathbb{R}$, we define the $\alpha$-wise components
@@
We henceforth refer to the two components with subscript ‘++’ as the plus components as they quantify how $F$ is, relative to $G$, shifted upwards and has an increased dispersion, respectively. Similarly, we refer to the terms with subscript ‘-’ as the minus components.
The decomposition terms in (7)–(8) attribute the overall difference between the central $\alpha$-interval endpoints of $F$ and GG—which is captured by the integrand $\operatorname{AVM}_{\alpha}(F,G)$ in (5)—to the components by more intricate interval comparisons. In a nutshell, the shift components capture twice the distance that one of the $\alpha$-intervals needs to be moved to lie within the other, while the dispersion components measure by how much one needs to (de)compress the $\alpha$-interval of $G$ to have the same length as the $\alpha$-interval of $F$.
@@
If, on the other hand, the $G$-interval is wider than the $F$-interval, the difference in interval length is attributed to the minus dispersion component. Analogously, if the intervals are ordered differently, their difference in location is attributed to the minus shift component. On a technical level, these separations into plus and minus terms are achieved by the $\cdot$$_{+}$ operator in (7) and (8), respectively.
@@ 2.2 The $p$-Wasserstein distance
We next present a generalized decomposition for the $p$-th power of the $p$-Wasserstein distance in (2) for $p\in[1,\infty)$. Here, we directly state the decomposition in its integral-form and dispense with a disaggregated $\alpha$-wise treatment as in (6), which can be obtained by simply omitting the integrals in the following formulas.
@@
Remark 2.4.
Irpino and Verde, (2015) discuss an alternative decomposition of the squared 2-Wasserstein distance into a location, scale and a remaining shape component given by
$\operatorname{WD}_{2}(F,G)=\left(\int_{0}^{1}|F^{-1}(\alpha)-G^{-1}(\alpha)|^{2}\text{d}\alpha\right)=\underbrace{(\mu_{F}-\mu_{G})^{2}}_{\text{location}}+\underbrace{(\sigma_{F}-\sigma_{G})^{2}}_{\text{scale}}+\underbrace{2\sigma_{F}\sigma_{G}(1-\rho_{F,G}),}_{\text{shape}}$ (13)
where $\mu_{F},\mu_{G}$ and $\sigma_{F},\sigma_{G}$ denote the respective means and standard deviations, while $\rho_{F,G}$ is the Pearson correlation of the points in a quantile-quantile plot of $F$ and $G$. In contrast to our quantile-based decomposition, the decomposition in (13) draws on the respective moments of the distributions. It could further be refined to directed location and scale components as in our decomposition by considering the sign of the differences in means and standard deviations, respectively.
The decomposition in (13) shares some of the properties of our decomposition for the CD and AVM, such as being shift invariant (i.e., simple shifts only affect the location component; c.f. Proposition 3.2) and compatibility with location-scale families (c.f. Proposition 3.4, which however does not apply to our $\operatorname{WD}_{p}$-decompositions for $p>1$). Furthermore, (13) is compatible with the usual stochastic and dispersive orders and is clearly linked to weak stochastic and dispersive orders based on ordering by means and standard deviations (cf. Section 4). The moment-based decomposition, however, is limited to the squared 2-WD and does not apply to the AVM and CD.
Remark 2.5.
While the $\operatorname{WD}_{p}$ decomposition from Proposition 2.3 holds for any finite $p\in[1,\infty)$, it is not obvious how this can be generalized to $p=\infty$, where
$\lim_{p\rightarrow\infty}\operatorname{WD}_{p}^{1/p}(F,G)=\sup_{\alpha\in(0,1)}|F^{-1}(\alpha)-G^{-1}(\alpha)|.$
In fact, our $\operatorname{WD}_{p}$ decomposition operates on the $p$-th power of the $p$-Wasserstein distance, and there is no canonical way of transposing it to the ordinary Wasserstein distance $\operatorname{WD}_{p}^{1/p}$. One possibility that achieves a decomposition of $\operatorname{WD}_{p}^{1/p}$ for fixed $p<\infty$ is to simply rescale the $\operatorname{WD}_{p}$ decomposition terms with $\frac{\operatorname{WD}_{p}^{1/p}}{\operatorname{WD}_{p}}$, i.e., for any $\operatorname{Comp}\in\{\operatorname{Shift}_{\pm},\operatorname{Disp}_{\pm}\}$ we set $\operatorname{Comp}^{\operatorname{WD}_{p}^{1/p}}=\frac{\operatorname{WD}_{p}^{1/p}}{\operatorname{WD}_{p}}\operatorname{Comp}^{\operatorname{WD}_{p}}$, which clearly yields rescaled components that sum to the $p$-th root $\operatorname{WD}_{p}^{1/p}$.
Then, taking the limit as $p\to\infty$ yields a decomposition for $\operatorname{WD}_{\infty}$. We suspect that these proportions will converge under fairly general conditions if the $\infty$-Wasserstein distance is finite. For example, Theorem 3.11 below (together with Proposition 3.3) implies that this is the case for the shift components when comparing symmetric distributions.
@@ 2.3 The Cramér distance
$\begin{aligned} \operatorname{Shift}_{+}^{\operatorname{CD}}(F,G):=\frac{1}{2}\int_{0}^{1}\int_{0}^{1}&\Big{[}\min\Big{\{}F^{-1}\left(\tfrac{1+\alpha}{2}\right)-G^{-1}\left(\tfrac{1+\beta}{2}\right),F^{-1}\left(\tfrac{1-\alpha}{2}\right)-G^{-1}\left(\tfrac{1-\beta}{2}\right)\Big{\}}\Big{]}_{+}\\ &+\left[F^{-1}\left(\tfrac{1-\alpha}{2}\right)-G^{-1}\left(\tfrac{1+\beta}{2}\right)\right]_{+}\,\mathrm{d}\alpha\,\mathrm{d}\beta,\end{aligned}$ (16)
@@ 3.1 Basic properties
For any $p\in[1,\infty)$ and $\operatorname{D}\in\{\operatorname{AVM},\operatorname{WD}_{p},\operatorname{CD}\}$, we have
@@
Hence, the properties that are outlined for the plus components (with subscript ‘++’) also apply to the minus counterparts (with subscript ‘-’) by symmetry.
@@
Let $F$ and $G$ be symmetric distributions with central medians $m_{F}$ and $m_{G}$ and $\operatorname{D}\in\{\operatorname{AVM},\operatorname{WD}_{p},\operatorname{CD}\}$, $p\in[1,\infty)$.
@@
    For any $\operatorname{D}\in\{\operatorname{AVM},\operatorname{WD}_{p},\operatorname{CD}\}$ with $p\in[1,\infty)$, it holds that
@@
    The central medians of $F$ and $G$ are $m_{F}=s_{F}m_{H}+\ell_{F}$ and $m_{G}=s_{G}m_{H}+\ell_{G}$, respectively, where $m_{H}$ denotes the central median of $H$. For any $p\in[1,\infty)$, if $m_{F}\leq m_{G}$, then $\operatorname{Shift}_{+}^{\operatorname{WD}_{p}}(F,G)=0$. Moreover, if the median of $H$ is unique,
@@
Remark 3.6.
The properties established in this section refer to the (non-)positivity of certain components, and it generally does not seem possible to make statements of a more quantitative nature. As an exception, for the dispersion components of the AVM, it holds that
$\operatorname{Disp}_{+}^{\text{AVM}}(F,G)-\operatorname{Disp}_{-}^{\text{AVM}}(F,G)=D_{F}-D_{G},$
where $D_{F}$ and $D_{G}$ are the mean absolute deviations around the medians of $F$ and $G$.
@@
We finally consider the case of normal distributions, which, arguably, form the most prominent location-scale family. Let $F=\mathcal{N}(\mu_{F},\sigma_{F}^{2})$ and $G=\mathcal{N}(\mu_{G},\sigma_{G}^{2})$ be normal distributions with means $\mu_{F}$ and $\mu_{G}$ and variances $\sigma_{F}^{2}$ and $\sigma_{G}^{2}$, respectively. In what follows, we use the shorthand notations $\widetilde{\mu}=|\mu_{F}-\mu_{G}|$, $\widetilde{\sigma}=|\sigma_{F}-\sigma_{G}|$ and $\widetilde{\rho}=\sqrt{\sigma_{F}^{2}+\sigma_{G}^{2}}$. We further denote the standard normal density and distribution function by $\phi(\cdot)$ and $\Phi(\cdot)$, respectively. Here, we provide closed-form expressions for the decompositions of the $\operatorname{AVM}$ and the $\operatorname{CD}$. More involved formulas for the $\operatorname{WD}_{p}$ with $p\in\mathbb{N}$ are given in Supplement S2.1.
@@
[…]
@@
$).\displaystyle\operatorname{Shift}_{+}^{\operatorname{CD}}(F,G)=\widetilde{\mu}\big{(}1-2\Phi(-\widetilde{\mu}/\widetilde{\rho})\big{)}-2\tau\big{(}\phi(0)-\phi(\widetilde{\mu}/\widetilde{\rho}$$\big{)}$.
@@ 3.2 Agreement and differences across distance measures
For any two distributions $F$ and $G$, and $p\in[1,\infty)$, it holds that
@@
For any two distributions $F$ and $G$, and $p\in[1,\infty)$, it holds that
@@
Let $F$ and $G$ be symmetric distributions, $F\neq G$, and $p,q\in[1,\infty)$. If $q>p$, it holds that
@@ 4 Compatibility with stochastic order relations
Dispersive preorder $F\geq_{\mathrm{D}}G$ $F^{-1}(\tau)-F^{-1}(\xi)\;\geq\;G^{-1}(\tau)-G^{-1}(\xi)\quad\forall\,0<\xi<\tau<1$
Weak dispersive preorder $F\geq_{\mathrm{wD}}G$ $F^{-1}(\tau)-F^{-1}(1-\tau)\;\geq\;G^{-1}(\tau)-G^{-1}(1-\tau)\quad\forall\,0.5<\tau<1$
@@
Usual stochastic partial order $F\geq_{\mathrm{S}}G$ $\min\big{\{}F^{-1}(\tau)-G^{-1}(\tau),\,F^{-1}(1-\tau)-G^{-1}(1-\tau)\big{\}}\;\geq\;0\quad\forall\,0.5<\tau<1$
Weak stochastic preorder (cond.) $F\geq_{\mathrm{wS}}G$ $\max\{F^{-1}(\tau)-G^{-1}(\xi),\,F^{-1}(1-\tau)-G^{-1}(1-\xi)\}\;\geq\;0\quad\forall\,0.5<\tau,\xi<1$
Relaxed stochastic preorder (cond.) $F\geq_{\mathrm{rS}}G$ $\max\{F^{-1}(\tau)-G^{-1}(\tau),\,F^{-1}(1-\tau)-G^{-1}(1-\tau)\}\;\geq\;0\quad\forall\,0.5<\tau<1$
Strong stochastic strict partial order $F>_{\mathrm{sS}}G$ $F\geq_{\mathrm{S}}G$ and $\exists\tau\in[0,1]:\;\;\min\{F(\tau)-G(\tau),\,F(1-\tau)-G(1-\tau)\}\;>\;0$
@@ 4.1 Connections to the dispersive order
The distribution $F$ is said to be larger than $G$ in dispersive order, $F\geq_{\text{D}}G$, if for all $0<\xi<\tau<1$,
@@
The dispersive order turns out to be too strong for a logical equivalence with zero dispersion components to hold. To establish such an equivalence, we consider a weaker order relation. The distribution $F$ is said to be larger than $G$ in weak dispersive order, $F\geq_{\mathrm{wD}}G$, if for all $0.5<\tau<1$,
@@
Let $F$ and $G$ be probability distributions with continuous quantile functions $F^{-1}$ and $G^{-1}$, respectively. Then, for any $\operatorname{D}\in\{\operatorname{AVM},\operatorname{WD}_{p},\operatorname{CD}\}$, $p\in[1,\infty)$, we have that
@@
Subject to the regularity condition of continuous quantile functions, which is further discussed below, part (a) of the theorem establishes an equivalence between a weak dispersive ordering of the distributions and a zero dispersion component in our decompositions. As shown in part (b), the equivalence extends naturally to an equivalence between a corresponding strict ordering and a unique nonzero dispersion component. Notably, the equivalences and implications in Theorem 4.1 hold for all considered divergence measures $\operatorname{D}\in\{\operatorname{AVM},\operatorname{WD}_{p},\operatorname{CD}\}$, $p\in[1,\infty)$, by the equivalences from Proposition 3.8.
@@
The issue encountered in Example 4.3 could be avoided by slightly adapting the definition of the weak dispersive order in (25) to require for any $0.5\leq\tau<1$ that
@@ 4.2 Connections to the usual stochastic order
Figure 8: Connections between the stochastic order relations introduced in Section 4.2 and the shift components of the $\operatorname{CD}$ and $\operatorname{WD}_{p}$ divergence measures (for any $p\in[1,\infty)$).
@@
The distribution $F$ is said to be larger than $G$ in the usual stochastic order, $F\geq_{\mathrm{S}}G$, if $F(x)\leq G(x)$ for all $x\in\mathbb{R}$, or, equivalently, if $F^{-1}(\tau)\geq G^{-1}(\tau)$ for all $0<\tau<1$. The usual stochastic order is a partial order, and its definition can be reformulated as
$\min\big{\{}F^{-1}(\tau)-G^{-1}(\tau),\,F^{-1}(1-\tau)-G^{-1}(1-\tau)\big{\}}\;\geq\;0$ (26)
for all $0.5<\tau<1$. The term used for this unconventional characterization arises in the shift components (see (7) with $\tau=\frac{1+\alpha}{2}$), thereby serving as a natural starting point to investigate the connection.